4 is 25% of x. Solve for x.
I would appreciate any help i can get on this! Thanks in advance!

Answer by rapaljer(4671)

Translate the sentence:
4 is 25% of x
4 = .25*x
4=1/4x

Multiply both sides by 4:
4*4=4*1/4*x
16=x

R^2